# Late-Time Evolution of Magnetized Disks in Tidal Disruption Events

### v1

@@ 2 Magnetized TDE Disks
where $T_{\rm eff}=T\left(\frac{4}{3\kappa_{\rm es}\Sigma}\right)^{1/4}$ is the effective temperature [^4] For simplicity we ignore the color corrections to emission that cause quasi-thermal accretion disk spectra to deviate from perfect blackbodies (Shimura & Takahara, 1995). While these corrections can cause $\sim 50\%$ differences between color temperature and effective temperature in X-ray bands, the effect is much smaller at longer wavelengths due to the increased absorption opacity in cooler parts of the disk (Davis & El-Abd, 2019)..
The effective viscosity is an approximation to the transport of angular momentum by magnetized turbulence (driven by the MRI), and when the pressure in the disk is dominated by magnetic fields, it becomes equal to:
$\nu_{\rm mag}$ $=\nu_{0,\rm{m}}R^{5/7}\Sigma^{2/7}$ (5)
$\nu_{0,\rm{m}}$ $=\left[\frac{27}{32\sigma_{\rm SB}}\alpha^{8}\kappa_{\rm{es}}GM_{\bullet}\left(\frac{k_{\rm B}}{\mu m_{\rm p}}\right)\right]^{1/7}.$
Because $\nu$ is a function of both $R$ and $\Sigma$ the evolution equation Eq. 1 becomes a nonlinear diffusion equation. This equation can be separated into a trivial time equation and a nonlinear ordinary differential equation for the radius (see Appendix A, where we show for the first time that a broad class of $\alpha$-viscosity laws can reduce the disk diffusion partial differential equation to the Emden-Fowler ordinary differential equation). However, there is currently no general analytical solution for this equation, necessitating the use of numerical methods.
2.1 Numerical Solution
In this paper, we integrate Eq. 1 numerically. We set the inner boundary to be the innermost stable circular orbit (ISCO), which for a non-spinning BH equals to $R_{\rm in}=6r_{\rm g}$, where $r_{\rm g}=GM_{\bullet}/c^{2}$ is the gravitational radius of the central BH and $c$ is the speed of light. We assume $\Sigma=0$ at the inner boundary, implies that the viscous torque there is zero, meaning that all the matter that reaches the ISCO is accreted along with its angular momentum. The outer boundary is always chosen to be $R_{\rm out}\gg R_{\rm in}$ such that the surface density $\Sigma$ vanishes at some radius smaller than $R_{\rm out}$ at all times.
In our fiducial models, we use a model in which, initially, half of the mass from the disrupted star has escaped to infinity, while the other half has already circularized to form a ring-like disk at (Strubbe & Quataert, 2009):
$r_{\rm c}=\frac{1+e_{\star}}{\beta}r_{\rm t}=2r_{\rm t}$ (6)
where $e_{\star}$ is the orbital eccentricity of the initially approaching star, and $\beta$ is the ratio of the tidal disruption radius $r_{\rm t}$ to the pericenter distance, and in the final equality we chose $e_{\star}=1$ and $\beta=1$. The tidal radius is (Rees, 1988):
$r_{\rm t}$ $=\left(\frac{M_{\bullet}}{M_{\star}}\right)^{1/3}R_{\star}$ (7)
$\sim 16\left(\frac{M_{\bullet}}{10^{6}M_{\odot}}\right)^{-2/3}\left(\frac{M_{\star}}{0.1M_{\odot}}\right)^{-1/3}\left(\frac{R_{\star}}{0.16R_{\odot}}\right)r_{\rm g}$
where $M_{\star}$ and $R_{\star}$ are the mass and radius of the disrupted star, respectively. We assume a lower-main-sequence star such that $R_{\star}=(M_{\star}/M_{\odot})^{0.8}R_{\odot}$ (Kippenhahn et al., 2012).
@@ 2.1 Numerical Solution
As a second scenario, instead of starting with a Gaussian initial condition, we also explore Eq. 1 with a source term that reflect the slow accumulation of the bound debris into the accretion disk. Initially, the surface density is a small value, and over time mass is added to the disk around $r_{\rm c}$ with a Gaussian distribution. The late-time rate of mass fallback onto the disk is (Stone et al., 2013):
@@
Figure 2: A comparison between the Gaussian initial condition (solid) to the source function (dashed). The top panel shows the surface density as a function of radius at different times, while the bottom panel presents the light curves at different wavelengths (NUV: $\lambda=250\text{nm}$, FUV: $\lambda=150\text{nm}$, and X-ray: $h\nu=300\text{eV}$). The masses of the SMBH and the disrupted star are $M_{\bullet}=10^{6}M_{\odot}$ and $M_{\star}=0.3M_{\odot}$ respectively, and $\alpha=0.1$.
The source function method is probably a more realistic solution than the initial Gaussian distribution at early times, although both approaches gloss over the considerable complexity (Hayasaki et al., 2013; Shiokawa et al., 2015; Bonnerot & Stone, 2021) of the circularization process. However, in this paper, we are interested in the late stages of the TDE, typically years after disruption. After such a long period, the initial conditions in our 1D models are effectively forgotten (Cannizzo et al., 1990), and there is not a big difference between the two prescriptions for mass injection. In Fig. 2, we compare the two methods, showing the surface density as a function of radius at different times, as well as the light curves. Although the surface densities differ quite a bit initially, they converge to the same solution over time. The UV light curves are also similar at late times, and while the X-ray luminosity is different, it decays rapidly and becomes effectively undetectable (Lodato & Rossi, 2011). Therefore, for simplicity, we use only the Gaussian initial condition in this paper unless specified otherwise.
@@ 2.2 Self-Similar solution
Although the evolution equation is a nonlinear equation without an exact analytic solution, in the case of highly magnetized accretion disks, where the viscosity [^5] Self-similar solutions can also exist for other types of strongly magnetized disk models, e.g. those where angular momentum transport is dominated by wind losses (Tamilan et al., 2025). follows a power law function of $\Sigma$ and $R$ (see Eq. 5), a self-similar solution for $\Sigma$ does exist (Lin & Pringle, 1987; Cannizzo et al., 1990; Pringle, 1991). In this self-similar solution, the inner boundary is at $R_{\rm{in}}=0$, and there are two types of solutions: (i) a constant disk mass with increasingly disk angular momentum, and (ii) a disk with zero torque at the origin. In this paper, we consider the latter solution, and for our model of a magnetized accretion disk the similarity solution is given by:
@@
×$\times\left[1-\frac{1}{52}\left(\frac{R}{R_{0,\rm S}}\right)^{13/9}\left(\frac{3\nu_{0,\rm S}t}{4R_{0,\rm S}^{2}}\right)^{-13/18}\right]^{7/2}$
@@
The self-similar solution is not the exact solution to Eq. 1, but it closely approximates the behavior of the system at late times. In Fig. 3, we compare the self-similar solution to the numerical solution for $M_{\star}=0.3M_{\odot}$. For the comparison, we need to determine $R_{0,\rm S}$, $\Sigma_{0,\rm S}$ and $\nu_{0,\rm S}$. By choosing $\nu_{0,\rm S}=\nu_{\rm mag}(R_{0,\rm S},\Sigma_{0,\rm S})$, we can rewrite Eq. 11 with only one free parameter:
@@
×$\times\left[1-\frac{1}{52}R^{13/9}\left(\frac{3\nu_{0,\rm S}t}{4}\right)^{-13/18}\right]^{7/2}$
where $\nu_{0,\rm{m}}$ is calculated at Eq. 5. To determine the remaining free parameter, $\nu_{0,\rm S}$, we calculated the outer radius of the numerical disk, i.e. the radius where the surface density drops to the numerical floor $\epsilon$, as a function of time (between $1$-$10$ years). We then fixed $\nu_{0,\rm S}$ to best fit the outer radius at these times. Fig. 3 shows that the self-similar solution closely resemble the numerical solution, with some differences: at large radii, the numerical solution decays to the numerical floor $\epsilon$, while the self-similar solution drops to zero. Additionally, differences in the inner boundary conditions result in different surface density profiles at small radii. In Appendix B, we provide a fitting function for $\nu_{0,\rm S}$, and present the relative errors between the light curves obtained from the numerical and the self-similar solutions.
@@ 3 Comparison to Linear Viscosity
In time-dependent 1D accretion disk models where radiation pressure dominates, the disk exhibits viscous and thermal instabilities. However, as described in Sec. 1, observations of many such astrophysical systems reveal disks that remain stable. A common theoretical solution is to modify the viscosity parametrization to a linear form (e.g. Mummery & Balbus, 2019):
@@
This linear viscosity approach is popular because it stabilizes the disk and is also theoretically simple, offering an analytical solution for Eq. 1, see for example Metzger et al. (2012). However, this choice of effective viscosity is generally made for numerical convenience and lacks a solid physical basis. Therefore, we aim to compare 1D disks evolving under linear viscosity model with strongly magnetized $\alpha$-disks. To do this, we explore various values of the power $\mu_{\rm l}$, and different methods for determining the normalization factor $\nu_{0,\rm{l}}$.
@@
In Fig. 4, we compare the surface densities $\Sigma$ between the linear viscosity and the magnetized disk models for $\mu_{\rm l}=\{3/2,0,-3/2\}$ at different times. The initial and boundary conditions are identical across all cases, and because we are particularly interested in late times, we determined the normalization factor $\nu_{0,\rm l}$ by equating the outer radii of the disks at 1 year (we explore different choices in Appendix C). In both the linear and the magnetized viscosity cases, the disk is stable, and the material that was once concentrated in a ring will gradually spread. Additionally, the surface densities at the outer radii are similar due to our choice of the normalization factor, but at small radii they are quite different. The best match between the surface density profiles occurs for $\mu_{\rm l}=0$, with worse agreement for other values of $\mu_{\rm l}$.
@@
However, we cannot observe the surface density directly, so we also need to look at the light curves. Therefore, in Fig. 5 we compare the light curves at three different wavelengths (near-UV, far-UV and X-ray), for both linear and $\alpha$-viscosity models. Surprisingly, the light curves for magnetic viscosity and linear viscosity with $\mu_{\rm l}=0$ looks very similar, even for the X-ray band that originates from the inner radii of the disk (Lodato & Rossi, 2011); other choices of $\mu_{\rm l}$ produce much worse agreement.
@@
Figure 4: Surface densities plotted against radius at different times, for the magnetized viscosity (solid) and the linear viscosity (dashed). The power laws of the linear viscosity are $\mu_{\rm l}=\{3/2,0,-3/2\}$ from the top panel to the bottom panel. The normalization factor of the linear parametrization $\nu_{0,\rm l}$ is determined by equating the outer radius of the two disks after 1 year. The choice of $M_{\bullet}$, $M_{\star}$, and $\alpha$ is the same as in Fig. 3. The best match between the surface densities occurs for $\mu_{\rm l}=0$.
@@
Figure 5: Light curves for near-UV, far-UV and X-ray bands (top panel to bottom panel) with Gaussian initial conditions for magnetized viscosity solutions (solid red) and linear viscosity solutions (dashed), with $\mu_{\rm l}=\{3/2,0,-3/2\}$ (black, blue, brown). The normalization factor of the linear parametrization $\nu_{0,\rm l}$ is determined by equating the outer radius of the disk after 1 year. The choice of $M_{\bullet}$, $M_{\star}$, and $\alpha$ is the same as in Fig. 3. The light curves of the linear and magnetized viscosity models appear very similar for $\mu_{\rm l}=0$, but quite distinct for other values of $\mu_{\rm l}$.
@@
To understand the similarity of the light curves between the magnetized $\alpha$-disk and the linear viscosity solutions with $\mu_{\rm l}=0$, we can check the time evolution of the self-similar solutions. From Eq. 11 the time evolution of the surface density of the magnetized disk is $\Sigma_{\rm m,S}\propto t^{-35/36}$, while the time evolution of the linear viscosity disk is $\Sigma_{\rm l,S}\propto t^{-(5-2\mu_{\rm l})/(4-2\mu_{\rm l})}$ (Metzger et al., 2008). For $\mu_{\rm l}=0$, the power laws are ultimately different, as can be seen in Fig. 4. However, in steady state, the luminosity $L\propto\dot{M}\propto\nu\Sigma$ where $\dot{M}$ is the mass accretion rate. When we substitute the self-similar solutions into the luminosity, we find:
@@
which for $\mu_{\rm l}=0$ reduce to precisely the same power-laws. This means that analytic Green’s function solutions with $\mu_{\rm l}=0$ can be used as a decent approximation to a magnetized accretion disk with $\alpha$-viscosity; these solutions take the form of Bessel functions in Newtonian gravity (e.g. Metzger et al., 2008) or hypergeometric functions in relativistic gravity (Mummery, 2023).
@@ 4 Observability of Late-Time TDE Disk Emission
Another insight from Fig. 5 is that TDEs remain UV-bright for many years. In this section, we aim to determine for just how long a TDE remains observable. Fig. 6 shows light curves at a UV wavelength of $\lambda=250\rm{nm}$ over thousands of years for different SMBH masses, and the Shakura-Sunyaev $\alpha$ parameters. As the SMBH mass increases, the TDE is brighter (Mummery et al., 2024), and for larger values of $\alpha$, the light curves evolve faster. To study the UV light curve, we define $R_{\rm UV}$ as the disk annulus where the temperature corresponds to the UV wavelength: $T_{\rm eff}(R_{\rm UV})=T_{\rm UV}=1.2\times 10^{4}\left(\frac{\lambda}{250\rm{nm}}\right)\rm{K}$. At this radius, the luminosity is $L_{\rm UV}\propto R^{2}_{\rm UV}\sigma_{\rm SB}T^{4}_{\rm UV}\propto\left(\nu\Sigma\right)^{2/3}$. Substituting the self-similar solution for a magnetized disk, the UV luminosity evolves over time as:
@@
Eventually, however, the TDE disk will turn off in the UV, either when the $R_{\rm UV}$ annulus becomes effectively optically thin or when it shrinks to the ISCO (in reality there will still be some UV emission after one of these criteria is satisfied, but it will become dimmer much more rapidly than in Eq. 15). The disk can be treated as a blackbody when the effective optical depth $\tau_{\rm eff}=\sqrt{3\kappa_{\rm abs}(\kappa_{\rm es}+\kappa_{\rm abs})}\Sigma>1$ for a given disk annulus, where $\kappa_{\rm abs}$ is the absorption opacity, approximated here by Kramers’ opacity $\kappa_{\rm abs}\sim\kappa_{\rm Kr}=1\times 10^{24}\left(\frac{\rho}{\rm g/cm^{3}}\right)\left(\frac{T}{\rm K}\right)^{-7/2}\rm{cm^{2}\,g^{-1}}$. As shown in Fig. 6, at larger SMBH masses, the disk will become optically thin more quickly, and $R_{\rm UV}$ will reach the ISCO sooner than is the case for smaller $M_{\bullet}$ values. Sometimes, the UV-emitting regions will become optically thin (in the effective opacity sense) before reaching the ISCO, the alternate ordering of events is also possible. Late-time TDEs are more likely to turn off in the UV due to optical depth effects when $\alpha$ is large and more likely to turn off due to $R_{\rm UV}\to R_{\rm ISCO}$ when $\alpha$ is small.

### v2

@@ 2.1 Numerical Solution
where $M_{\star}$ and $R_{\star}$ are the mass and radius of the disrupted star, respectively. We assume a lower-main-sequence star such that $R_{\star}=(M_{\star}/M_{\odot})^{0.8}R_{\odot}$ (Kippenhahn+2012).
@@
As a second scenario, instead of starting with a Gaussian initial condition, we also explore Eq.˜1 with a source term that reflect the slow accumulation of the bound debris into the accretion disk. Initially, the surface density is a small value, and over time mass is added to the disk around $r_{\rm c}$ with a Gaussian distribution. The late-time rate of mass fallback onto the disk is (Stone+2013):
@@
Figure 2: A comparison between the Gaussian initial condition (solid) to the source function (dashed). The top panel shows the surface density as a function of radius at different times, while the bottom panel presents the light curves at different wavelengths (NUV: $\lambda=250\text{nm}$, FUV: $\lambda=150\text{nm}$, and X-ray: ${\color[rgb]{1,0,0}\textbf{}}hf=300\text{eV}$). The masses of the SMBH and the disrupted star are $M_{\bullet}=10^{6}M_{\odot}$ and $M_{\star}=0.3M_{\odot}$ respectively, and $\alpha=0.1$.
The source function method is probably a more realistic solution than the initial Gaussian distribution at early times, although both approaches gloss over the considerable complexity (Hayasaki+2013; Shiokawa+2015; Bonnerot&Stone2021) of the circularization process. However, in this paper, we are interested in the late stages of the TDE, typically years after disruption. After such a long period, the initial conditions in our 1D models are effectively forgotten (Cannizzo+1990), and there is not a big difference between the two prescriptions for mass injection. In Fig.˜2, we compare the two methods, showing the surface density as a function of radius at different times, as well as the light curves. Although the surface densities differ quite a bit initially, they converge to the same solution over time. The UV light curves are also similar at late times, and while the X-ray luminosity is different, it decays rapidly and becomes effectively undetectable (Lodato&Rossi2011). Therefore, for simplicity, we use only the Gaussian initial condition in this paper unless specified otherwise.
@@ 2.2 Self-Similar solution
Although the evolution equation is a nonlinear equation without an exact analytic solution, in the case of highly magnetized accretion disks, where the viscosity [^5] Self-similar solutions can also exist for other types of strongly magnetized disk models, e.g. those where angular momentum transport is dominated by wind losses (Tamilan+2025). follows a power law function of $\Sigma$ and $R$ (see Eq.˜5), a self-similar solution for $\Sigma$ does exist (Lin&Pringle1987; Cannizzo+1990; Pringle1991). In this self-similar solution, the inner boundary is at $R_{\rm{in}}=0$, and there are two types of solutions: (i) a constant disk mass with increasingly disk angular momentum, and (ii) a disk with zero torque at the origin. In this paper, we consider the latter solution, and for our model of a magnetized accretion disk the similarity solution is given by:
@@
$\times\left[1-\frac{1}{52}\left(\frac{R}{R_{0,\rm S}}\right)^{13/9}\left(\frac{3\nu_{0,\rm S}t}{4R_{0,\rm S}^{2}}\right)^{-13/18}\right]^{7/2}$
@@
The self-similar solution is not the exact solution to Eq.˜1, but it closely approximates the behavior of the system at late times. In Fig.˜3, we compare the self-similar solution to the numerical solution for $M_{\star}=0.3M_{\odot}$. For the comparison, we need to determine $R_{0,\rm S}$, $\Sigma_{0,\rm S}$ and $\nu_{0,\rm S}$. By choosing $\nu_{0,\rm S}=\nu_{\rm mag}(R_{0,\rm S},\Sigma_{0,\rm S})$, we can rewrite Eq.˜11 with only one free parameter:
@@
$\times\left[1-\frac{1}{52}R^{13/9}\left(\frac{3\nu_{0,\rm S}t}{4}\right)^{-13/18}\right]^{7/2}$
where $\nu_{0,\rm{m}}$ is calculated at Eq.˜5. The outer radius of the disk in the self-similar solution is $R_{\rm out}=52^{9/13}\left(\frac{3\nu_{0,\rm S}t}{4}\right)^{1/2}\propto t^{1/2}$. To determine the remaining free parameter, $\nu_{0,\rm S}$, we calculated the outer radius of the numerical disk, i.e. the radius where the surface density drops to the numerical floor $\epsilon$, as a function of time (between 11-$10$ years). We then fixed $\nu_{0,\rm S}$ to best fit the outer radius at these times. Fig.˜3 shows that the self-similar solution closely resemble the numerical solution, with some differences: at large radii, the numerical solution decays to the numerical floor $\epsilon$, while the self-similar solution drops to zero. Additionally, differences in the inner boundary conditions result in different surface density profiles at small radii. In Appendix B, we provide a fitting function for $\nu_{0,\rm S}$, and present the relative errors between the light curves obtained from the numerical and the self-similar solutions.
@@ 3 Comparison to Linear Viscosity
In time-dependent 1D accretion disk models where radiation pressure dominates, the disk exhibits viscous and thermal instabilities. However, as described in Sec. 1, observations of many such astrophysical systems reveal disks that remain stable. A common theoretical solution is to modify the viscosity parametrization to a linear form (e.g. Mummery2019):
@@
This linear viscosity approach is popular because it stabilizes the disk and is also theoretically simple, offering an analytical solution for Eq.˜1, see for example Metzger+2012. However, this choice of effective viscosity is generally made for numerical convenience and lacks a solid physical basis. Therefore, we aim to compare 1D disks evolving under linear viscosity model with strongly magnetized $\alpha$-disks. To do this, we explore various values of the power $\mu_{\rm l}$, and different methods for determining the normalization factor $\nu_{0,\rm{l}}$.
One value of $\mu_{\rm l}$ that may be relevant at late times can be derived from the self-similar solution. From Eq.˜11, we find that $\Sigma\propto R^{-5/9}$, and substituting this into Eq.˜5 gives $\nu_{\rm mag}\propto R^{\mu_{\rm l}}$, where $\mu_{\rm l}=5/9$. Throughout this section, we also explore other values for the power law index: $\mu_{\rm l}=\{3/2,0,-3/2\}$.
@@
In Fig.˜4, we compare the surface densities $\Sigma$ between the linear viscosity and the magnetized disk models at different times. The initial and boundary conditions are identical across all cases, and because we are particularly interested in late times, we determined the normalization factor $\nu_{0,\rm l}$ by equating the outer radii of the disks at 1 year (we explore different choices in Appendix C). In both the linear and the magnetized viscosity cases, the disk is stable, and the material that was once concentrated in a ring will gradually spread. Additionally, the surface densities at the outer radii are similar due to our choice of the normalization factor, but at small radii they can differ significantly. The best match between the surface density profiles occurs for $\mu_{\rm l}=5/9$ at late-times. However, $\mu_{\rm l}=0$ provides the best agreement with the time evolution in the outer regions of the disk. Other values of $\mu_{\rm l}$ show worse agreement with the behavior of the magnetized disk.
@@
However, we cannot observe the surface density directly, so we also need to look at the light curves. Therefore, in Fig.˜5 we compare the light curves at three different wavelengths (near-UV, far-UV and X-ray), for both linear and $\alpha$-viscosity models. Surprisingly, the light curves for magnetic viscosity and linear viscosity with $\mu_{\rm l}=0$ looks very similar, even for the X-ray band that originates from the inner radii of the disk (Lodato&Rossi2011); other choices of $\mu_{\rm l}$ produce much worse agreement.
@@
Figure 4: Surface densities plotted against radius at different times, for the magnetized viscosity (solid) and the linear viscosity (dashed). The power laws of the linear viscosity are $\mu_{\rm l}=\{3/2,5/9,0,-3/2\}$ from the top panel to the bottom panel. The normalization factor of the linear parametrization $\nu_{0,\rm l}$ is determined by equating the outer radius of the two disks after 1 year. The choice of $_{\bullet}, MM_{\star}$, and $\alpha$ is the same as in Fig.˜3. The best match between the surface densities occurs for $\mu_{\rm l}=5/9$ in the inner regions of the disk and $\mu_{\rm l}=0$ in the outer regions.
@@
Figure 5: Light curves for near-UV, far-UV and X-ray bands (top panel to bottom panel) with Gaussian initial conditions for magnetized viscosity solutions (solid red) and linear viscosity solutions (dashed), with $\mu_{\rm l}=\{3/2,5/9,0,-3/2\}$ (green, black, blue, brown). The normalization factor of the linear parametrization $\nu_{0,\rm l}$ is determined by equating the outer radius of the disk after 1 year. The choice of $_{\bullet}, MM_{\star}$, and $\alpha$ is the same as in Fig.˜3. The light curves of the linear and magnetized viscosity models appear very similar for $\mu_{\rm l}=0$, but quite distinct for other values of $\mu_{\rm l}$.
@@
To understand the similarity of the light curves between the magnetized $\alpha$-disk and the linear viscosity solutions with $\mu_{\rm l}=0$, we can check the time evolution of the self-similar solutions. From Eq.˜11 the time evolution of the surface density of the magnetized disk is $\Sigma_{\rm m,S}\propto t^{-35/36}$, while the time evolution of the linear viscosity disk is $\Sigma_{\rm l,S}\propto t^{-(5-2\mu_{\rm l})/(4-2\mu_{\rm l})}$ (Metzger+2008). For $\mu_{\rm l}=0$, the power laws are ultimately different, as can be seen in Fig.˜4. However, in steady state, the luminosity $L\propto\dot{M}\propto\nu\Sigma$ where $\dot{M}$ is the mass accretion rate. When we substitute the self-similar solutions into the luminosity, we find:
@@
which for $\mu_{\rm l}=0$ reduce to precisely the same power-laws. This means that analytic Green’s function solutions with $\mu_{\rm l}=0$ can be used as a decent approximation to a magnetized accretion disk with $\alpha$-viscosity; these solutions take the form of Bessel functions in Newtonian gravity (e.g. Metzger+2008) or hypergeometric functions in relativistic gravity (Mummery2023).
@@ 4 Observability of Late-Time TDE Disk Emission
Another insight from Fig.˜5 is that TDEs remain UV-bright for many years. In this section, we aim to determine for just how long a TDE remains observable. Fig.˜6 shows light curves at a UV wavelength of $\lambda=250\rm{nm}$ over thousands of years for different SMBH masses, and the Shakura-Sunyaev $\alpha$ parameters. As the SMBH mass increases, the TDE is brighter (Mummery+2024), and for larger values of $\alpha$, the light curves evolve faster. To study the UV light curve, we define $R_{\rm UV}$ as the disk annulus where the temperature corresponds to the UV wavelength: $T_{\rm eff}(R_{\rm UV})=T_{\rm UV}=1.2\times 10^{4}\left(\frac{\lambda}{250\rm{nm}}\right)\rm{K}$. At this radius, the luminosity is $L_{\rm UV}\propto R^{2}_{\rm UV}\sigma_{\rm SB}T^{4}_{\rm UV}\propto\left(\nu\Sigma\right)^{2/3}$. Substituting the self-similar solution for a magnetized disk, the UV luminosity evolves over time as:
@@
Eventually, however, the TDE disk will turn off in the UV, either when the $R_{\rm UV}$ annulus becomes effectively optically thin or when it shrinks to the ISCO (in reality there will still be some UV emission after one of these criteria is satisfied, but it will become dimmer much more rapidly than in Eq. 15). The disk can be treated as a blackbody when the effective optical depth $\tau_{\rm eff}=\sqrt{3\kappa_{\rm abs}(\kappa_{\rm es}+\kappa_{\rm abs})}\Sigma>1$ for a given disk annulus, where $\kappa_{\rm abs}$ is the absorption opacity, approximated here by Kramers’ opacity $\kappa_{\rm abs}\sim\kappa_{\rm Kr}=1\times 10^{24}\left(\frac{\rho}{\rm g/cm^{3}}\right)\left(\frac{T}{\rm K}\right)^{-7/2}\rm{cm^{2}\,g^{-1}}$. As shown in Fig.˜6, at larger SMBH masses, the disk will become optically thin more quickly, and $R_{\rm UV}$ will reach the ISCO sooner than is the case for smaller $M_{\bullet}$ values. Sometimes, the UV-emitting regions will become optically thin (in the effective opacity sense) before reaching the ISCO, the alternate ordering of events is also possible. Late-time TDEs are more likely to turn off in the UV due to optical depth effects when $\alpha$ is large and more likely to turn off due to $R_{\rm UV}\to R_{\rm ISCO}$ when $\alpha$ is small.